SW&IT Meeting Minutes 2018-07-02
Attendees: Andrew, David, Doug, Igor, Kellee, Matthieu
, Petr, Stephen, Xianyu, Yang
News from management
Previous Action Items
- Git & releases - Thursday 28th June
Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.
This week in github
- Moved ARGOS wiki out of sandbox and published
- Built ARGOS from GitHub
- ARGOS docker and TAN NICE RPMS
- ARGOS swing arm and calibration unit in OSS
- Try to run ARGOS from Docker containers
- Ingest handover documents into CAN
- Ketiv pre-support
- Passwords, etc. that they really should have had
- Access setup and verification (Annoyingly, they are "frozen" at Teamviewer v.12, necessitating that we also do not update)
- Chasing managers for IssueTrak email list info
- Started looking into GitHub web redirect with Stephen
- Monitoring Vault upgrade (happening now)
- More manager chasing
- Backup testing
- Pointing Discussions & historical email transfer to Yang
- Modeling/analysis work for IE/CA and Pointing in prep for J.Hill mtg
- Follow-up from J.Hill meeting
- Completed the Binocular Planner, rev 1.0.0.
- It is ready for team test at this time.
- Demo of Planner for MB, ME, IS
- The way I wrote the planner, only exposure steps have offsets.
- Non exposure steps use the observation target for their sky position
- For Observations that start with a nonzero offset, this may place the preset at an unexpected location. Whether this is a bug or feature may be instrument-specific
- IT6544 OT Rules are incomplete CLOSED - The rules will probably be forever incomplete, as more are added with new instruments coming online. Not a bug.
- IT7183 LBT 2018AV1.0.0 Unable to Generate LBC Scripts CLOSED. I have had no further reports of issues since I fixed this in early April. Bugs fixed.
- No action items this week.
- Fixed the LBC power GUI that had a "refresh" bug between blue filter wheel 1 and 2 because of cable swaps awhile ago.
- Wrote up more LBC minicom notes in the wiki troubleshooting, including some details for a filter wheel test.
- Created a new wiki page for the kernel update to the OVMS racktangle: OVMSRackTangleKernelUpdate
- git discussions and experimentation for TCS upgrades to GCS.
- Started testing the GCS updates for the OAC HK upgrade.
- Added a new error case to OAC for when the startAGW fails due to the new 8051 tripping a threshold. Built this new version in the 32bit environment on the mountain.
- Update wiki for leap seconds detail changes.
- Get GCS OAC HK updates into git branch.
- half-day holiday for 4th of July
- Finished GCS image flip
- Worked on IT3xxx (Time To Live, looks I have the cause of wrong calculations)
- Worked and hopefully finished X flip for WFS Zernikes (IT7317), encountered problem during build (which I have to revisit)
- Looked on leap second issue, discovered reflective memory issue (IT7320)
- closed/put to monitor some older issues
- M3 summer shutdown activities
- OVMS SD card activities
- Time To Limits bugfixing
- Preparing 2018B release (Time To Limits, GCS integrals, ECS - changes in reflective memory)
- Build ARGOS docker image using cisrv01.tucson.lbto.org
- ARGOS TAN RPMs build (on CentOS 6.9,7) , but do not work with cisrv01.tucson.lbto.org
- SOUL jumbo frames
- Jonathan Crass, connecting network devices.
- Continue to study the AO supervisor code and the IT issue #7261 (memory problem with WfsArbitrator)
- Working with Doug to transition and learn the pointing model construction and improvement.
- Working on pointing model related wiki pages.
- Working on my own branch of AO supervisor code.